A department store manager has monitored the number of complaints received per week about poor service. The probabilities for numbers of complaints in a week, established by this review, are shown in the table. Number of complaints 0 1 2 3 4 5 Probability 0.17 0.30 0.33 0.07 0.05 0.08 What is the mean of complaints received per week?
Answer #1

Given that,

X 0 1 2 3 4 5
P(X) 0.17 0.30 0.33 0.07 0.05 0.08

We know formula for mean is as follows:

Mean(\mu) = \sum x\times p(x)

\therefore Mean(\mu) = (0\times0.17)+(1\times0.3)+(2\times0.33)+(3\times0.07) +(4\times0.05)+(5\times0.08)

\therefore Mean(\mu) = (0)+(0.30)+(0.66)+(0.21) +(0.20)+(0.40)

\therefore Mean(\mu) = 1.77

Hence, the mean of complaints received per week= 1.77